OpenStudy (anonymous):

5x^2-7x=4 find -13f(7)-6f(-10)

OpenStudy (anonymous):

first lets us know what is f(x)?

OpenStudy (anonymous):

f(x) is 5x^2-7x=4

OpenStudy (anonymous):

^2 means to the second power

OpenStudy (anonymous):

f(x, y) = x 3 + xy + y 3 + 1 f(0, 0) = 1 f(1/3, 1/3) = 32/27 f(−1/3, −1/3) = 28/27 0 0 = Df = h ∂f ∂x ∂f ∂y i = 3x 2 + y x + 3y 2

OpenStudy (anonymous):

is the answer corect

OpenStudy (anonymous):

helpingtutors

OpenStudy (anonymous):

i dont understand the answer, the questions says, if f(x)=5x^2-7x=4, find -13f(7)-6f(-10)

OpenStudy (anonymous):

Does it perhaps say f(x)=5x^2-7x+4? The equals there makes no sense.

OpenStudy (anonymous):

then it says find the foll owing values: -13f(7)-6f(-10)

OpenStudy (anonymous):

Ok, then. \[f(x)=5x^2-7x+4\] \[f(7)=5*7^2-7*7+4\]\[f(7)=5*49-49+4\]\[f(7)=200\] \[f(-10)=5*-10^2-7*(-10)+4\]\[f(7)=5*100+70+4\]\[f(7)=574\] \[-13f(7)-6f(-10)=-13*200-6*574 = -6 044\]

OpenStudy (anonymous):

Those 7s should be -10s in the second paragraph

OpenStudy (anonymous):

so the second answer is wrong?

OpenStudy (anonymous):

No, I just forgot to put -10s in the brackets (I copied and pasted)

OpenStudy (anonymous):

i'm sorry but how did you get -10?

OpenStudy (anonymous):

You were asked to find -13f(7)-6f(-10) In the first paragraph thing, I worked out f(7) and in the second I did f(-10)

OpenStudy (anonymous):

yes